# Towards Unbiased and Robust Spatio-Temporal Scene Graph Generation and Anticipation

### v1

@@ Abstract
Spatio-Temporal Scene Graphs (STSGs) provide a concise and expressive representation of dynamic scenes by modelling objects and their evolving relationships over time. However, real-world visual relationships often exhibit a long-tailed distribution, causing existing methods for tasks like Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) to produce biased scene graphs. To this end, we propose ImparTail, a novel training framework that leverages curriculum learning and loss masking to mitigate bias in the generation and anticipation of spatio-temporal scene graphs. Our approach gradually decreases the dominance of the head relationship classes during training and focuses more on tail classes, leading to more balanced training. Furthermore, we introduce two new tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—designed to evaluate the robustness of STSG models against distribution shifts. Extensive experiments on the Action Genome dataset demonstrate that our framework significantly enhances the unbiased performance and robustness of STSG models compared to existing methods.
Figure 1: (a) Long Tailed Distribution. Predicates in Spatio-Temporal Scene Graph (STSG) datasets exhibit a long-tailed distribution; one such example is the Action Genome [17] dataset, whose distribution is described at the top right. (b) Tasks. We focus on two STSG tasks, including Video Scene Graph Generation (VidSGG) on the left and Scene Graph Anticipation (SGA) on the right. VidSGG entails the identification of fine-grained relationships between the objects observed in the video, such as (Person, looking_at, Paper Notebook) and (Person, not_looking_at, Paper Notebook) in respective frames to the left. (c) Conventional Learning. Due to the inherent long-tailed distribution of these datasets, models learnt using the conventional approaches focus more on the head classes and perform poorly on the tail classes as illustrated using the prediction scores of STTran [7] on contacting and attention relationships (refer middle row). (d) Unbiased Learning. To alleviate the dominance of head classes during training, in unbiased learning, we focus more on the tail classes, ensuring that the learnt models exhibit significantly better performance in predicting both head and tail classes (refer bottom row).
@@ 1 Introduction
As AI systems evolve to interact with complex, real-world environments, accurate, adaptive modeling becomes essential. In particular, fields such as autonomous driving [23], human-robot interaction [29], and security surveillance [49] require predictive understanding that goes beyond conventional computer vision. Spatio-Temporal Scene Graphs (STSGs) represent a powerful approach that can address this challenge; at a high level, STSGs are structured graphs where nodes correspond to objects and edges capture the evolving relationships between them over time [17, 23]. Unlike static scene graphs, STSGs offer nuanced video understanding, enabling autonomous systems to interpret interactions in real-time, anticipate changes, and make informed decisions [41].
@@
Yet, constructing effective STSGs that mirror the intricacies of the dynamic, ever-changing real world remains an unsolved problem. In particular, real-world relationships follow a highly imbalanced, long-tailed distribution, with a handful of head classes appearing frequently [44, 2]. In contrast, the vast majority of so-called tail classes are rare but essential for a detailed understanding. This imbalance often skews the learned models towards common relationships, yielding an incomplete, incorrect, unsafe, and distorted interpretation of the scene [42]. Addressing this challenge is vital: without a nuanced, detailed, and balanced representation, AI systems risk overlooking the subtle yet critical interpretations that define complex scenes.
@@
To tackle this challenge, we introduce ImparTail, an innovative framework designed to overcome the pervasive long-tail distribution bias in STSG generation. Via a novel curriculum learning strategy [12], ImparTail shifts attention from common head classes to tail classes, promoting a balanced representation that accurately captures all relationship types, not just the dominant ones. Additionally, ImparTail employs loss masking techniques [27, 2] that amplify the influence of rare classes during training, ensuring a more equitable learning process. This controlled shift towards tail classes allows models to address data distribution disparities without compromising accuracy on head classes.
@@
Real-world data poses an additional challenge beyond bias: distributional shifts. Factors such as lighting, occlusions, and background changes often degrade model performance, affecting the reliability of scene graphs in real-world applications [54, 40]. Conventional models struggle to generalize across these shifts, limiting their effectiveness in unpredictable conditions. To address this, we introduce two new evaluation tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—that simulate realistic distributional shifts to assess model resilience under deployment-like conditions.
We rigorously assess the performance of ImparTail on the Action Genome dataset [17], a key benchmark for video understanding. Our results show that ImparTail achieves high mean recall scores. Additionally, ImparTail exhibits robust performance amid distributional shifts, demonstrating that it is well-suited for building robust real-world systems. Thus, this paper makes the following contributions:
@@
  • ImparTail. We introduce a novel, effective, unbiased learning framework for generating and anticipating Spatio-Temporal Scene Graphs by leveraging Partial Stochastic Gradient Descent with Curriculum Learning [12, 28]. Our method is especially suited to address the challenges posed by the long-tailed distribution of real-world relationships.
@@ 2 Related Work
Tasks. Image Scene Graph Generation (ImgSGG) focuses on representing static visual data—such as 2D and 3D images—as spatial graphs where objects are nodes and their relationships are edges. The field gained significant traction with the foundational Visual Genome project [24], which advanced research in 2D ImgSGG. Building upon this work, Kim et al. [23] extended the task to static 3D scene data by incorporating both RGB and depth information. Object interactions over time provide richer contextual information for dynamic visual content like videos. Transforming this content into structured Spatio-Temporal Scene Graphs (STSGs)—with nodes representing objects and edges capturing temporal relationships—is known as Video Scene Graph Generation (VidSGG). Research in VidSGG has concentrated on improving representation learning through advanced object-centric architectures such as STTran [7] and RelFormer [39]. Shifting the focus from the identification and generation of scene graphs, Peddi et al. [35] recently introduced the Scene Graph Anticipation (SGA) task, which aims to predict STSGs for future frames. Alongside these developments, foundation models have propelled advancements in various ImgSGG task variants [6, 22, 55, 53, 26].
Unbiased Learning. TEMPURA [30] and FlCoDe[21] address the challenges posed by long-tailed datasets, such as those found in Action Genome [17] and VidVRD [38] and propose methods for unbiased VidSGG. Specifically, FloCoDe [21] mitigates bias by emphasizing temporal consistency and correcting the imbalanced distribution of visual relationships. Similarly, TEMPURA [30] addresses biases in relationship prediction with memory-guided training to generate balanced relationship representations and applies a Gaussian Mixture Model to reduce predictive uncertainty. Note. To the best of our knowledge, we are the first to systematically investigate model biases in the SGA task and assess the robustness of both VidSGG and SGA models.
@@
Curriculum Learning (CL) is a training methodology that structures training by presenting simpler examples first and progressively introducing more complex ones. This approach aims to enhance learning efficiency by aligning the difficulty of training data with the model’s learning capacity at each stage [25, 11, 12, 13]. Despite its potential benefits, implementing CL presents significant challenges. A primary obstacle is the non-trivial task of distinguishing between easy and hard training samples. This differentiation often requires additional mechanisms, such as auxiliary neural networks acting as teachers or specialized algorithms. Difficulty measures can be predefined based on certain heuristics [16] or learned automatically during the training process [25, 18, 46, 19, 36, 12, 28]. Alongside difficulty assessment, a scheduling strategy is essential to determine when and how to introduce more challenging data [11].
Figure 2: Overview of ImparTail (a) Pipeline. The forward pass of ImparTail begins with an ORPU, where initial object proposals are generated for each observed frame. These object representations are then fed to STPUs designed to construct spatio-temporal context-aware relationship representations of interacting objects. ImparTail applied to both tasks VidSGG and SGA remains mostly the same, with an additional LDPU unit added for SGA to anticipate relationship representations for future frames. These observed (for VidSGG)/ anticipated (for SGA) relationship representations are then decoded to construct STSGs. (b) Conventional Loss. Previous approaches estimated loss for all relationship predicates (head and tail classes). (c) Masked Loss. With the inherent long-tailed nature of the STSG datasets, conventional training results in biased VidSGG and SGA models. Thus, to de-bias the training and learn an unbiased model, in ImparTail, instead of estimating loss for all relationship predicates, we estimate a masked loss, where we selectively mask the labels corresponding to dominant head classes and void their contribution in learning. (d) Curriculum Learning. In ImparTail, we introduce a curriculum-based approach for masking relationship predicate labels during training. At each iteration, we adjust the selection of masked predicates to balance the class distribution progressively. As illustrated, initially, the model trains on the original, long-tailed distribution. As training advances, we systematically mask predicate labels from the head classes, gradually shifting the distribution toward uniformity.
@@ 3 Notation & Background
We borrow the notation from SceneSayer[35] and briefly present it below. Let $\textstyle\mathcal{D}$ correspond to the ground truth annotations of an STSG dataset. Consider a video segment $\textstyle V_{1}^{T}$ from a video $\textstyle V$, modelled as a sequence of frames at discrete time steps $\textstyle t=1,2,\ldots,\bar{T}$:
$\textstyle V_{1}^{T}=\{I^{1},I^{2},\ldots,I^{T}\}.$
@@
Here, $\textstyle T$ denotes the total number of frames in the segment, which may vary between videos. In each frame $\textstyle I^{t}$, we represent the scene using a scene graph that captures the visual objects and their pairwise relationships. Let $\textstyle O^{t}=\{o_{1}^{t},o_{2}^{t},\ldots,o_{N(t)}^{t}\}$ be the set of observed objects in frame $\textstyle I^{t}$, where $\textstyle N(t)$ is the number of objects in that frame. Each object $\textstyle o_{k}^{t}$ is characterized by: A bounding box $\textstyle\mathbf{b}_{k}^{t}\in[0,1]^{4}$, representing its spatial location and A category label $\textstyle\mathbf{c}_{k}^{t}\in\mathcal{C}$, where $\textstyle\mathcal{C}$ is the set of all object categories. To describe relationships between objects, we define $\textstyle\mathcal{P}$ as the set of all predicate classes representing possible spatio-temporal relationships. For any pair of objects $\textstyle(o_{i}^{t},o_{j}^{t})$, multiple predicates $\textstyle\{p_{ijk}^{t}\}$ may be associated, where each predicate $\textstyle p_{ijk}^{t}\in\mathcal{P}$. A relationship instance is then defined as the triplet:
$\textstyle r_{ijk}^{t}=(o_{i}^{t},p_{ijk}^{t},o_{j}^{t}).$
Thus, the scene graph $\textstyle\mathcal{G}^{t}$ for frame $\textstyle t$ consists of all such relationship triplets:
$\textstyle\mathcal{G}^{t}=\{r_{ijk}^{t}\}_{i,j,k}.$
@@
For each observed object $\textstyle o_{i}^{t}$ and each object pair $\textstyle(o_{i}^{t},o_{j}^{t})$, we define probability distributions over object categories and predicate classes, respectively: $\textstyle\hat{\mathbf{c}}_{i}^{t}\in[0,1]^{|\mathcal{C}|}$ is the probability distribution over object categories for $\textstyle o_{i}^{t}$ and $\textstyle\hat{\mathbf{p}}_{ij}^{t}\in[0,1]^{|\mathcal{P}|}$ is the probability distribution over predicate classes for the pair $\textstyle(o_{i}^{t},o_{j}^{t})$. These distributions are normalized, satisfying:
$\textstyle\sum_{k}\hat{c}_{ik}^{t}=1,\quad\sum_{k}\hat{p}_{ijk}^{t}=1.$
@@
We distinguish and clarify the objectives of Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) as follows: (a) VidSGG. The primary goal of VidSGG is to construct a sequence of scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=1}^{T}$ corresponding to the observed video segment $\textstyle V_{1}^{T}=\{I^{t}\}_{t=1}^{T}$. This involves: Identifying objects $\textstyle\{o_{k}^{t}\}_{k=1}^{N(t)}$ within each frame $\textstyle I^{t}$ and Determining all possible pairwise relationships $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ among the detected objects. (b) SGA. The main objective of SGA is to generate scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=T+1}^{T+H}$ for future frames $\textstyle V_{T+1}^{T+H}=\{I^{t}\}_{t=T+1}^{T+H}$ of the video, using information from the observed segment $\textstyle V_{1}^{T}$. Here, $\textstyle H$ denotes the anticipation horizon [^1] Following the definition and assumptions from SceneSayer [35], we assume the continuity of observed objects in future frames—that is, $\textstyle\{o_{i}^{t}\}_{i=1}^{N(t)}=\{o_{i}^{T}\}_{i=1}^{N(T)}$ for all $\textstyle t>T$—and focus on predicting the evolution of their relationships $\textstyle\{\{r_{ijk}^{t}\}_{t=T+1}^{T+H}\}_{i,j,k}$..
Graph Building Strategies
Building upon the literature in Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA), we implement three distinct strategies for constructing scene graphs:
  1. 1.
@@
    With Constraint Strategy: This strategy imposes a unique interaction constraint between every pair of objects within a scene. Specifically, for any two objects $\textstyle(o_{i}^{t},o_{j}^{t})$, there exists exactly one relationship predicate $\textstyle p_{ij}^{t}$ (for each relationship category) that describes their interaction. We include these relationship triplets $\textstyle\{r_{ij}^{t}\}_{ij}$ in the scene graph $\textstyle\mathcal{G}^{t}$.
  2. 2.
    No Constraint Strategy: This approach allows for a more expressive graph structure by permitting multiple relationship predicates between any pair of interacting objects $\textstyle(o_{i}^{t},o_{j}^{t})$. We incorporate all predicted relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ into the scene graph $\textstyle\mathcal{G}^{t}$.
  3. 3.
    Semi-Constraint Strategy: This strategy strikes a balance by permitting multiple relationships between object pairs $\textstyle(o_{i}^{t},o_{j}^{t})$ only if their confidence scores exceed a predefined threshold. Specifically, we add relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ to the scene graph $\textstyle\mathcal{G}^{t}$ when the confidence score satisfies $\textstyle\hat{p}_{ijk}^{t}>\theta_{k}$.

### v2

@@ Abstract
Spatio-Temporal Scene Graphs (STSGs) provide a concise and expressive representation of dynamic scenes by modeling objects and their evolving relationships over time. However, real-world visual relationships often exhibit a long-tailed distribution, causing existing methods for tasks like Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) to produce biased scene graphs. To this end, we propose ImparTail, a novel training framework that leverages loss masking and curriculum learning to mitigate bias in the generation and anticipation of spatio-temporal scene graphs. Unlike prior methods that add extra architectural components to learn unbiased estimators, we propose an impartial training objective that reduces the dominance of head classes during learning and focuses on underrepresented tail relationships. Our curriculum-driven mask generation strategy further empowers the model to adaptively adjust its bias mitigation strategy over time, enabling more balanced and robust estimations. To thoroughly assess performance under various distribution shifts, we also introduce two new tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—offering a challenging benchmark for evaluating the resilience of STSG models. Extensive experiments on the Action Genome dataset demonstrate the superior unbiased performance and robustness of our method compared to existing baselines.
@@ 1 Introduction
Spatio-Temporal Scene Graphs (STSGs) are structured graphs where nodes correspond to objects and edges capture the evolving relationships between them over time [19, 25]. Unlike static scene graphs, STSGs offer nuanced video understanding, enabling autonomous systems to interpret interactions in real-time, anticipate changes, and make informed decisions [44]. Yet, constructing effective STSGs that mirror the intricacies of the dynamic, ever-changing real world remains an unsolved problem. In particular, real-world relationships follow a highly imbalanced, long-tailed distribution, with a handful of head classes appearing frequently [47, 2]. In contrast, the vast majority of so-called tail classes are rare but essential for a detailed understanding. This imbalance often skews the learned models towards common relationships, yielding an incomplete, incorrect, unsafe, and distorted interpretation of the scene [45]. Addressing this challenge is vital: without a nuanced, detailed, and balanced representation, AI systems risk overlooking the subtle yet critical interpretations that define complex scenes.
Figure 1: Overview. Row-1: Existing training pipelines in the literature for VidSGG/SGA tasks. Row-2: Prior unbiased learning work, exemplified by the supplementary architectural modules and loss functions. Row-3 (ImparTail): A framework through which any prior object-centric, representation learning–based VidSGG/SGA method can be adapted to learn corresponding unbiased estimator.
@@
To tackle this challenge, we introduce ImparTail, a novel framework designed to overcome long-tail distribution bias in STSG generation. ImparTail shifts attention from common head classes to tail classes, promoting a balanced representation that accurately captures all relationship types, not just the dominant ones. It employs a novel loss masking technique (similar in spirit to [30, 2]) that amplifies the influence of rare classes during training, ensuring an equitable learning process. This controlled shift towards tail classes allows models to address data distribution disparities without significantly compromising performance on head classes.
Figure 2: (a) Long Tailed Distribution. Predicates in Spatio-Temporal Scene Graph (STSG) datasets exhibit a long-tailed distribution; one such example is the Action Genome [19] dataset, whose distribution is described at the top left. (b) Tasks. We focus on two STSG tasks, including Video Scene Graph Generation (VidSGG) on the left and Scene Graph Anticipation (SGA) on the right. VidSGG entails the identification of fine-grained relationships between the objects observed in the video, such as (Person, looking_at, Paper Notebook) and (Person, not_looking_at, Paper Notebook) in respective frames to the left. SGA aims to anticipate the evolution of these relationships to (Person, touching, Cup), and eventually, (Person, drinking_from, Cup) [37]. (c) Conventional Learning. Due to the inherent long-tailed distribution of these datasets, models learnt using the conventional approaches focus more on the head classes and perform poorly on the tail classes as illustrated using the prediction scores of STTran [7] on contacting and attention relationships (refer middle row). (d) Unbiased Learning. To alleviate the dominance of head classes during training, in unbiased learning, we focus more on the tail classes, ensuring that the learnt models exhibit significantly better performance in predicting both head and tail classes (refer bottom row).
@@
Real-world data poses an additional challenge beyond bias: distributional shifts. Factors such as lighting, occlusions, and background changes often degrade model performance, affecting the reliability of scene graphs in real-world applications [57, 43]. Conventional models struggle to generalize across these shifts, limiting their effectiveness in unpredictable conditions. To address this, we introduce two new evaluation tasks—Robust Spatio-Temporal Scene Graph Generation and Robust Scene Graph Anticipation—that simulate realistic distributional shifts to assess model resilience under deployment-like conditions. Thus, this paper makes the following contributions [^1] Project: https://rohithpeddi.github.io/#/impartail:
@@
  • ImparTail. We introduce a novel, effective, unbiased learning framework for generating and anticipating Spatio-Temporal Scene Graphs by leveraging Curriculum-Guided Masked Training with Partial Gradients [14, 31]. Our method is especially suited to address the challenges posed by the long-tailed distribution of real-world relationships.
@@ 2 Related Work
Tasks. Image Scene Graph Generation (ImgSGG) focuses on representing static visual data—such as 2D and 3D images—as spatial graphs where objects are nodes and their relationships are edges. The field gained significant traction with the foundational Visual Genome project [26]. Building upon this work, Kim et al. [25] extended the task to static 3D scene data by incorporating both RGB and depth information. Object interactions over time provide richer contextual information for dynamic visual content like videos. Transforming this content into structured Spatio-Temporal Scene Graphs (STSGs)—with nodes representing objects and edges capturing temporal relationships—is known as Video Scene Graph Generation (VidSGG). Research in VidSGG has concentrated on improving representation learning through advanced object-centric architectures such as STTran [7] and RelFormer [42]. Shifting the focus from the identification and generation of scene graphs, Peddi et al. [37] recently introduced the Scene Graph Anticipation (SGA) task, which aims to predict STSGs for future frames.
Unbiased Learning. TEMPURA [32] and FlCoDe[23] address the challenges posed by long-tailed datasets, such as those found in Action Genome [19] and VidVRD [41] and propose methods for unbiased VidSGG [^2] To the best of our knowledge, we are the first to investigate biases in the SGA task and assess the robustness of both VidSGG and SGA models.. Specifically, FloCoDe [23] mitigates bias by emphasizing temporal consistency and correcting the imbalanced distribution of visual relationships. Similarly, TEMPURA [32] addresses biases in relationship prediction with memory-guided training to generate balanced relationship representations and applies a Gaussian Mixture Model to reduce predictive uncertainty.
@@
Curriculum Learning (CL) is a training methodology that structures training by presenting simpler examples first and progressively introducing more complex ones. This approach aims to enhance learning efficiency by aligning the difficulty of training data with the model’s learning capacity at each stage [27, 13, 14, 15]. Despite its potential benefits, implementing CL presents significant challenges. A primary obstacle is the non-trivial task of distinguishing between easy and hard training samples. Difficulty measures can be predefined based on certain heuristics [18] or learned automatically during the training process [27, 20, 49, 21, 38, 14, 31]. Alongside difficulty assessment, a scheduling strategy is essential to determine when and how to introduce more challenging data [13].
Figure 3: Overview of ImparTail (a) Pipeline. The forward pass of ImparTail begins with an ORPU, where initial object proposals are generated for each observed frame. These object representations are then fed to STPUs designed to construct spatio-temporal context-aware relationship representations of interacting objects. ImparTail applied to both tasks VidSGG and SGA remains mostly the same, with an additional LDPU unit added for SGA to anticipate relationship representations for future frames. These observed (for VidSGG)/ anticipated (for SGA) relationship representations are then decoded to construct STSGs. (b) Conventional Training. Previous approaches estimated loss for all relationship predicates (head and tail classes). (c) Masked Training. With the inherent long-tailed nature of the STSG datasets, conventional training results in biased VidSGG and SGA models. Thus, to de-bias the training and learn an unbiased model, in ImparTail, instead of estimating loss for all relationship predicates, we estimate a masked loss, where we selectively mask the labels corresponding to dominant head classes and void their contribution in learning. (d) Curriculum-Guided Mask Generation. In ImparTail, we introduce a curriculum-based approach for masking relationship predicate labels during training. At each iteration, we adjust the selection of masked predicates to balance the class distribution progressively. As illustrated, initially, the model trains on the original, long-tailed distribution. As training advances, we systematically mask predicate labels from the head classes, gradually shifting the distribution toward uniformity.
@@ 3 Notation & Background
We borrow the notation from SceneSayer[37] and briefly present it below. Let $\textstyle\mathcal{D}$ correspond to the ground truth annotations of an STSG dataset. Consider a video segment $\textstyle V_{1}^{T}$ from a video $\textstyle V$, modeled as a sequence of frames at discrete time steps $\textstyle t=1,2,\ldots,T$: $\textstyle V_{1}^{T}=\{I^{1},I^{2},\ldots,I^{T}\}$ where $\textstyle T$ denotes the total number of frames in the segment. In each frame $\textstyle I^{t}$, we represent the scene using a scene graph that captures the visual objects and their pairwise relationships. Let $\textstyle O^{t}=\{o_{1}^{t},o_{2}^{t},\ldots,o_{N(t)}^{t}\}$ be the set of observed objects in frame $\textstyle I^{t}$, where $\textstyle N(t)$ is the number of objects in that frame. Each object $\textstyle o_{k}^{t}$ is characterized by: A bounding box $\textstyle\mathbf{b}_{k}^{t}\in[0,1]^{4}$, representing its spatial location and A category label $\textstyle\mathbf{c}_{k}^{t}\in\mathcal{C}$, where $\textstyle\mathcal{C}$ is the set of object categories.
@@
To describe relationships between objects, we define $\textstyle\mathcal{P}$ as the set of all predicate classes representing possible spatio-temporal relationships. For any pair of objects $\textstyle(o_{i}^{t},o_{j}^{t})$, multiple predicates $\textstyle\{p_{ijk}^{t}\}$ may be associated, where each predicate $\textstyle p_{ijk}^{t}\in\mathcal{P}$. A relationship instance is then defined as the triplet: $\textstyle r_{ijk}^{t}=(o_{i}^{t},p_{ijk}^{t},o_{j}^{t})$. Thus, the scene graph $\textstyle\mathcal{G}^{t}$ for frame $\textstyle t$ consists of all such relationship triplets: $\textstyle\mathcal{G}^{t}=\{r_{ijk}^{t}\}_{i,j,k}$. For each observed object $\textstyle o_{i}^{t}$ and each object pair $\textstyle(o_{i}^{t},o_{j}^{t})$, we define probability distributions over object categories and predicate classes, respectively: $\textstyle\hat{\mathbf{c}}_{i}^{t}\in[0,1]^{|\mathcal{C}|}$ is the probability distribution over object categories for $\textstyle o_{i}^{t}$ and $\textstyle\hat{\mathbf{p}}_{ij}^{t}\in[0,1]^{|\mathcal{P}|}$ is the probability distribution over predicate classes for the pair $\textstyle(o_{i}^{t},o_{j}^{t})$. These distributions are normalized, satisfying: $\textstyle\sum_{k}\hat{c}_{ik}^{t}=1,\quad\sum_{k}\hat{p}_{ijk}^{t}=1.$
@@
We distinguish and clarify the objectives of Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA) as follows: (a) VidSGG. The primary goal of VidSGG is to construct a sequence of scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=1}^{T}$ corresponding to the observed video segment $\textstyle V_{1}^{T}=\{I^{t}\}_{t=1}^{T}$. This involves: Identifying objects $\textstyle\{o_{k}^{t}\}_{k=1}^{N(t)}$ within each frame $\textstyle I^{t}$ and Determining all possible pairwise relationships $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ among the detected objects. (b) SGA. The main objective of SGA is to generate scene graphs $\textstyle\{\mathcal{G}^{t}\}_{t=T+1}^{T+H}$ for future frames $\textstyle V_{T+1}^{T+H}=\{I^{t}\}_{t=T+1}^{T+H}$ of the video, using information from the observed segment $\textstyle V_{1}^{T}$. Here, $\textstyle H$ denotes the anticipation horizon [^3] Following the definition and assumptions from SceneSayer [37], we assume the continuity of observed objects in future frames..
@@
Graph Building Strategies. Following the literature in Video Scene Graph Generation (VidSGG) and Scene Graph Anticipation (SGA), we implement three distinct strategies for constructing scene graphs: (a) With Constraint Strategy: This strategy imposes a unique interaction constraint between every pair of objects within a scene. Specifically, for any two objects $\textstyle(o_{i}^{t},o_{j}^{t})$, there exists exactly one relationship predicate $\textstyle p_{ij}^{t}$ (for each relationship category) that describes their interaction. We include these relationship triplets $\textstyle\{r_{ij}^{t}\}_{ij}$ in the scene graph $\textstyle\mathcal{G}^{t}$. (b) No Constraint Strategy: This approach allows for a more expressive graph structure by permitting multiple relationship predicates between any pair of interacting objects $\textstyle(o_{i}^{t},o_{j}^{t})$. We incorporate all predicted relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ into the scene graph $\textstyle\mathcal{G}^{t}$. (c) Semi-Constraint Strategy: This strategy strikes a balance by permitting multiple relationships between object pairs $\textstyle(o_{i}^{t},o_{j}^{t})$ only if their confidence scores exceed a predefined threshold. Specifically, we add relationship triplets $\textstyle\{r_{ijk}^{t}\}_{i,j,k}$ to the scene graph $\textstyle\mathcal{G}^{t}$ when the confidence score satisfies $\textstyle\hat{p}_{ijk}^{t}>\theta_{k}$.